# Effect Size Estimation for Duration Recommendation in Online Experiments: Leveraging Hierarchical Models and Objective Utility Approaches

### v1

@@ Abstract
The selection of the assumed effect size (AES) critically determines the duration of an experiment, and hence its accuracy and efficiency. Traditionally, experimenters determine AES based on domain knowledge. However, this method becomes impractical for online experimentation services managing numerous experiments, and a more automated approach is hence of great demand. We initiate the study of data-driven AES selection in for online experimentation services by introducing two solutions. The first employs a three-layer Gaussian Mixture Model considering the heteroskedasticity across experiments, and it seeks to estimate the true expected effect size among positive experiments. The second method, grounded in utility theory, aims to determine the optimal effect size by striking a balance between the experiment’s cost and the precision of decision-making. Through comparisons with baseline methods using both simulated and real data, we showcase the superior performance of the proposed approaches.
1 Introduction
@@
Sample size determination (SSD) plays a pivotal role in online experiments, answering the critical question of "how long should an experiment run?" Richardson et al. (2022). In the spheres of online A/B testing and related classic application such as clinical trials, it is paramount to ascertain the minimal sample size during the experiment planning phase. If the sample size is too small, it may not accurately represent the whole population being studied, thereby introducing both bias in the results that limit drawing conclusions or generalizing the inferences. It can also compromise the statistical power of the experiment, leading to inaccuracies in detecting meaningful effects Ramsey and Schafer (2012); Lenth (2001). Conversely, if the sample size is selected as too large, it could unnecessarily extend the duration of the experiment and inflate the associated costs, such as the human and hardware resources and opportunity costs Wan et al. (2023). Therefore, selecting an appropriate sample size is imperative to maintain a balance between decision accuracy and resource utilization.
@@
Among the numerous methodologies for SSD in the literature  Kelley and Rausch (2006); Adcock (1997); Lindley (1997); Weiss (1997), one crucial step is to specify the assumed effect size (the absolute or percent lift from the treatment over the control). For example, it is a critical component in statistical power analysis used for SSD, in either the frequentist or the Bayesian setting  Du and Wang (2016). First of all, we need to distinguish the true effect size of an experiment and the assumed effect size (AES). The first is an objective unknown quantity, while the latter is a subjective manually specified number. In traditional power analysis problems (e.g., in clinical trials), AES mainly reflects (i) the experimenter’s expectation (e.g., what level of improvement would be regarded as acceptable) and (ii) the trade-off between the opportunity cost of running a longer experiment and the accuracy. AES does not necessarily relate to the true effect size (or its estimate). Moreover, using observed power based on the true effect size of an experiment to determine its sample size is inappropriate, due to its 1-1 mapping to p-value Hoenig and Heisey (2001). For example, in case where an experiment lacks a noticeable effect size, the resulting observed power tends to be small. Using this value in SSD can lead to overestimating the required experiment duration.
@@
2 Related Work
Our work is closely related to the SSD problem, for which numerous methodologies have been studied in the literature. For example, one approach involves selecting a sample size to achieve a narrow confidence interval for the standardized mean difference Kelley and Rausch (2006). Alternatively, sample sizes can be determined based on utility theory Adcock (1997); Lindley (1997) or to ensure the Bayes factor exceeds a predetermined value Weiss (1997). However, in all these works, the effect size is assumed as a hyper-parameter that has been pre-specified by experimenters, which is not practical for online experimentation services.
@@ 2 Related Work
3 Preliminaries and Notations
Online A/B experiments are widely used to compare customer responses between the existing feature A and the new feature B, guiding online companies’ feature launch decisions. To reduce the noise and established the causal relationship in online experiments, it is standard to track customers’ outcomes (e.g. ad clicks) after a customer is triggered based on some event (e.g., displaying ads). Let the duration of the experiment be defined in terms of a certain number of weeks, which helps mitigate the influence of daily and weekend fluctuations. To maintain the simplicity, we assume the experiments have the same duration $t$ and omit the time subscript $t$ in notations. Denote $n_{g}$ as the total number of customers who are triggered up to week $t$ in group $g\in\{T,C\}$ (T for the treatment and C for control). Denote sets of those customers who are triggered up to week $t$ as $I_{g}$. Denote $Y_{i}$ as the observed outcome for customer $i$ up to week $t$, and $\bar{Y}_{g}=\sum_{i\in I_{g}}\frac{1}{n_{g}}Y_{i}$ is the sample mean of responses for customers in group $g$. Assuming that $Y_{i}$ are independent and identically distributed (i.i.d.) with mean $u_{g}$ and finite variance $\sigma^{2}_{g}$, as $n_{.}\rightarrow\infty$, we can apply the central limit theorem (CLT) Casella and Berger (2021):
@@ 3 Preliminaries and Notations
$H_{0}\mathrel{\mathop{\ordinarycolon}}\mu\leq 0\quad H_{1}\mathrel{\mathop{\ordinarycolon}}\mu>0.$
@@
where $\alpha$ is Type I error. $\delta$ is AES. $z_{.}$ represents the standard normal quantile. $\Phi(.)$ stands for the cumulative distribution function of the standard normal distribution. To achieve a desired power of $b$ for the alternative hypothesis, considering for a given effect size $\delta>0$ and an allocation rate of $n_{T}/n_{C}=p$, the minimum sample size $n_{T},n_{C}$ are selected to satisfying $power(\delta,\sigma_{.},n_{.},\alpha)\geq b$. The required minimum duration can be determined through sample size prediction Richardson et al. (2022).
Standardized effect sizes, such as Cohen’s d Cohen (2013), are often used in meta-analysis across different experiments, serving to standardize variations in ATE magnitudes. If the estimators of standardized effect size follow an asymptotic normal distribution, all models presented in this paper can seamlessly accommodate the use of standardized effect sizes. Consequently, a dedicated discussion on this topic is omitted in this paper.
3.1 Pooled Effect Size
@@
The following model is used to estimate the AES Du and Wang (2016):
@@
In an online company setting, not every experiment yields a measurable effect size. This stems from the nature of online experiments with their short cycles, which was designed to encourage the exploration of innovative ideas. However, the drawback is that a significant number of ideas may not achieve statistical significance. We categorize experiments into three groups: 1) true positive, 2) true negative, and 3) flat experiments which have no significant effect. Using $d_{i}$ from all experiments to train the model (3.1) will lead to an underestimation of AES, consequently elongating the required duration. The guideline is to train the model (3.1) using experiments where we know there is an underlying true positive effect. While it might seem intuitive to use hypothesis testing to categorize experiments into significant positive, negative or flat ones, the existence of Type I and Type II errors poses a challenge in precisely determining which experiments truly have such effects.
4 Three-Layer Heteroscedastic GMM
@@
(c) The simulated true effect sizes under increased random errors, with three clusters corresponding to positive, flat, and negative.
@@
In section 3.1, we highlighted that the main challenge of the state-of-the-art method lies in the accurately identifying experiments with true positive effect. It is natural to consider the categorization of experiments as latent variables and use the Gaussian Mixture Model (GMM).
Motivation. We first illustrate our motivation of using GMM in Figure 1. Figure 0(a) presents the distribution of the observed effect sizes for certain outcome of interest over 3,300 experiment run within Amazon. Although at the first glance, this graph seems to support a normal distribution, we argue that it indeed illustrates the challenge of this problem. In Figure 0(b), we simulate the true effect sizes following a two-Layer GMM, and in Figure 0(c) we increases random errors to each of them to generate the observed effect size, which looks like a single mode Gaussian distribution. Therefore, under the assumption that there exists latent clusters of positive, flat and negative experiments, it is actually infeasible to identify those positive ones with simple rules as in Section 3; instead, a more principled statistical model as GMM should be used. The two-layer GMM didn’t account for the required heterogeneity in our setting (Section 3). Thus, we propose a three-layer heteroscedastic GMM.
@@
Latent variable $\bm{z}=\{z_{1},z_{2},\ldots,z_{m}\}$, $z_{i}\in\{1,\ldots,K\}$ represent the index class sampled from the categorical distribution parameterized by $\bm{\pi}=(\pi_{1},\ldots,\pi_{K})$. $\mu_{k},\tau_{k}$ and $\pi_{k}$ correspond to the mean, variance, and weight of the $k$-th Gaussian component. Denote $\bm{\theta}=\{\mu_{1},\ldots,\mu_{K},\tau_{1},\ldots,\tau_{K}\}$. We have
@@
where $\sigma_{i}$ are known and can be estimated using observed outcomes of the experiment $i$, e.g. Equation (1). We can further pre-specify the means of k = 2 cluster as zero, i.e. $\mu_{2}=0$.
@@
$=$ $\sum_{i=1}^{m}\mathbb{E}_{z|y}\bigg{[}log\bigg{(}p_{i}(d_{i},z_{i}|\bm{\theta},\bm{\pi})\bigg{)}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}$
$=$ $\sum_{i=1}^{m}\sum_{k=1}^{K}\mathbb{E}_{z|y}\bigg{[}{\mathbb{I}[z_{i}=k]}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}\times\bigg{[}log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))\bigg{]}$
@@
The EM algorithm is summarized in Algorithm 1. By setting $K=3$ and $\mu_{2}=0$, the mean for positive components $\mu_{1}$ serves as the estimate of AES. To avoid getting trapped in stationary pointWu (1983), we repeat the EM algorithm multiple times with randomly initialized starting points.
Singularity issues in Gaussian mixture model. When the covariance matrix is singular, the variances becomes zero, it leads to a spiky Gaussian component that "collapses" into a single point. There are several papers discussing how to address this issue. The first approach involves setting a lower bound on variance Hathaway (1985). The second approach introduces a penalty term in the log-likelihood function to prevent the variance of a specific component from becoming too small Chen et al. (2008); Ridolfi and Idier (2001). In addition, Chen et al. (2008); Chen (2017) introduce the conditions and provide proofs for the consistency of the MLE under finite Gaussian mixture model. Using the second approach, the marginal log-likelihood takes the following form:
$log(f(\bm{d}|\bm{\theta},\bm{\pi}))+log(p_{m}(\bm{\tau^{2}}))=\sum_{i=1}^{m}\sum_{k=1}^{K}[log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))]-\frac{1}{m}\sum_{k=1}^{K}(\frac{1}{\tau^{2}_{k}}+log(\tau^{2}_{k})),$ (10)
@@
5 Utility Theory: Bayesian Optimal Effect Size
In Section 4, we proposed a modeling approach that focuses on estimating the mean effect size. Such an approach is easy to explain to experimentation service users. However, in many real-world scenarios, users prioritize identifying the most suitable parameters for achieving maximum gains rather than just maximizing the accuracy of estimation. Motivated by the real needs, we propose a utility-maximization framework as our second approach.
@@ 5 Utility Theory: Bayesian Optimal Effect Size
$\operatorname*{arg\;max}_{\mu>0}\;\;\sum_{i=1}^{m}{\mathbb{E}}_{\delta_{i}\sim\mathcal{N}(d^{\prime}_{i},\sigma_{i}^{2})}\Bigg{[}\underbrace{-c_{i}\times(T_{i}(\mu)-1)}_{\text{Opportunity Cost}}+\;\underbrace{\big{(}\delta_{i}\times N_{i,1\mathrel{\mathop{\ordinarycolon}}T_{i}(\mu),T}\big{)}}_{\mathclap{\begin{subarray}{c}\text{Impact during the experiment}\\ \text{(relative to control)}\end{subarray}}}+\;\;\underbrace{u_{2}(\delta_{i},H,T_{i}(\mu))\mathbb{I}(\pi(\bm{Y}_{T_{i}(\mu),i})=1)}_{\text{Launch impact (relative to control)}}\Bigg{]}$ (13)
$=\operatorname*{arg\;max}_{\mu>0}\sum_{i=1}^{m}\Bigg{[}\underbrace{-c_{i}\times(T_{i}(\mu)-1)}_{\text{Opportunity Cost}}+\;\;\underbrace{\big{(}d^{\prime}_{i}\times N_{i,1\mathrel{\mathop{\ordinarycolon}}T_{i}(\mu),T}\big{)}}_{\mathclap{\begin{subarray}{c}\text{Impact during the experiment}\\ \text{(relative to control)}\end{subarray}}}+\;\;\;\underbrace{u_{2}(d^{\prime}_{i},H,T_{i}(\mu))\mathbb{I}(\pi(\bm{Y}_{T_{i}(\mu),i})=1)}_{\text{Launch impact (relative to control)}}\Bigg{]},$ (16)
@@
  • $N_{i,1\mathrel{\mathop{\ordinarycolon}}t,g}$ denotes the total number of customers in group $g\in\{T,C\}$ observed up to week $t$ for experiment $i$.
@@
  • $u_{2}(\delta,H,t)$ represents the impact of launching the treatment feature on the whole population at week $t$, given a pre-specified time horizon $H$. Typically, we study one year impact, i.e. $H=52$ weeks. One example of estimating one year launch impact is: $u_{2}(\delta,H,t)=\delta*(H-t)*(\sum_{g\in\{T,C\}}N_{i,1\mathrel{\mathop{\ordinarycolon}}t,g})$
@@
The equality in Equation (13) is due to the linearity of expectation when $u_{2}$ is a linear function in the ATE (which is true in our case).
Optimization. Equation (13) is a one-dimensional optimization problem, for which we can efficiently find a good solution. We use grid search in our prototype.
@@ 6 Experiments
6.1 Accuracy Comparison with Simulation
@@
Methods and Results. Table 1 compares the MSE and MAE for the pooled effect size (Pooled-MLE), standard two-layer GMM (Two-layer GMM) and the proposed three-layer heteroscedastic GMM (Three-layer GMM). For Pooled-MLE, the MLE of parameter $\mu_{0}$ is computed using the model in Equation (3.1) on all positive observed effect sizes $d_{i}$. Two-layer GMM uses Gaussian Mixture in sklearn package Pedregosa et al. (2011) with $K=3$. Three-layer GMM, the proposed method, uses the proposed EM-Algorithm 1 with the penalty term (10), $K=3$ and $\mu_{2}=0$. Each simulation uses $10$ different starting points to avoid local optimum (One of these starting points is initialized using the k-means clustering). Both Two-layer GMM and Three-layer GMM uses the estimated mean for the positive components as AES and $tolerance=10^{-3}$. Table 1 shows that Three-layer GMM performs slightly better than Two-layer GMM. The p-value from a two-sample t-test is 0.036, indicating the Three-layer GMM has significantly better accuracy. As expected, Pooled-MLE is underestimated. GMM outperforms Pooled-MLE notably in cases where we lack information about whether the experiment’s effect is a true positive or not. The Boxplot in Figure 3 illustrates a similar conclusion.
@@
6.2 Meta-Analysis with Real Experiments
@@
Dataset and setup. We collect a dataset of 3,300 historical experiments conducted in the past two years within Amazon from the United States, each having a duration of 4 weeks. $d_{i}$ represents the observed standardized effect size derived from a specific standardization formulation used in the company at week 4. Given the absence of the ground-truth effect size, we adopt a heuristic yet easy-to-explain approach that uses the observed effect size at the end of the 4 weeks as the empirical ground-truth effect size.
We set the maximum duration as $4$ weeks for the analysis. For each estimated AES obtained from different methods, we plug in them, along with the observed sample size and sample variance (or their predicted versions as described in Richardson et al. (2022)), into power Equation (2) to calculate the statistical power at week $i=1,2,3,4$. We define the recommended duration as the minimum number of weeks ($\leq 4$) required to attain a power of 80%. In Frequentist setting, we use the one-sided two-group Welch’s t-test Welch (1947) as decision policy $\pi(.)$. This implies that the decision policy $\pi(.)=1$ for launch if the p-value is less than the significance level of $0.05$ and the ATE is positive.
Table 3: Simulated data analysis results (Ground truth is 1).
@@
Methods. We compare the proposed estimated AESs from Three-Layer GMM and utility theory-based optimal effect size (Utility-maximization) with those from two-layer GMM and Pooled-MLE. The settings for Two-layer GMM, Three-layer GMM and Pooled_MLE are the same as Section 6.1. Utility-maximization uses grid-search to find the optimal effect size within the set $\{0.02\%,0.04\%,\ldots,2\%\}$.
@@
    The average utility in Equation (13) and its three components, including the opportunity cost, the impact during the experiment, and the launch impact.

### v2

@@ Abstract
The selection of the assumed effect size (AES) critically determines the duration of an experiment, and hence its accuracy and efficiency. Traditionally, experimenters determine AES based on domain knowledge. However, this method becomes impractical for online experimentation services managing numerous experiments, and a more automated approach is hence of great demand. We initiate the study of data-driven AES selection for online experimentation services by introducing two solutions. The first employs a three-layer Gaussian Mixture Model considering the heteroskedasticity across experiments, and it seeks to estimate the true expected effect size among positive experiments. The second method, grounded in utility theory, aims to determine the optimal effect size by striking a balance between the experiment’s cost and the precision of decision-making. Through comparisons with baseline methods using both simulated and real data, we showcase the superior performance of the proposed approaches.
Introduction
@@
Sample size determination (SSD) plays a pivotal role in online experiments, answering the critical question of ”how long should an experiment run?” (Richardson et al. 2022). In the spheres of online A/B testing and related classic application such as clinical trials, it is paramount to ascertain the minimal sample size during the experiment planning phase. If the sample size is too small, it may not accurately represent the whole population being studied, thereby introducing both bias in the results that limit drawing conclusions or generalizing the inferences. It can also compromise the statistical power of the experiment, leading to inaccuracies in detecting meaningful effects (Ramsey and Schafer 2012; Lenth 2001). Conversely, if the sample size is selected as too large, it could unnecessarily extend the duration of the experiment and inflate the associated costs, such as the human and hardware resources and opportunity costs (Wan et al. 2023). Therefore, selecting an appropriate sample size is imperative to maintain a balance between decision accuracy and resource utilization.
@@
Among the numerous methodologies for SSD in the literature  (Kelley and Rausch 2006; Adcock 1997; Lindley 1997; Weiss 1997), one crucial step is to specify the assumed effect size (the absolute or percent lift from the treatment over the control). For example, it is a critical component in statistical power analysis used for SSD, in either the frequentist or the Bayesian setting  (Du and Wang 2016). First of all, we need to distinguish the true effect size of an experiment and the assumed effect size (AES). The first is an objective unknown quantity, while the latter is a subjective manually specified number. In traditional power analysis problems (e.g., in clinical trials), AES mainly reflects (i) the experimenter’s expectation (e.g., what level of improvement would be regarded as acceptable) and (ii) the trade-off between the opportunity cost of running a longer experiment and the accuracy. AES does not necessarily relate to the true effect size (or its estimate). Moreover, using observed power based on the true effect size of an experiment to determine its sample size is inappropriate, due to its 1-1 mapping to p-value (Hoenig and Heisey 2001). For example, in case where an experiment lacks a noticeable effect size, the resulting observed power tends to be small. Using this value in SSD can lead to overestimating the required experiment duration.
@@
Related Work
Our work is closely related to the SSD problem, for which numerous methodologies have been studied in the literature. For example, one approach involves selecting a sample size to achieve a narrow confidence interval for the standardized mean difference (Kelley and Rausch 2006). Alternatively, sample sizes can be determined based on utility theory (Adcock 1997; Lindley 1997) or to ensure the Bayes factor exceeds a predetermined value (Weiss 1997). However, in all these works, the effect size is assumed as a hyper-parameter that has been pre-specified by experimenters, which is not practical for online experimentation services.
@@
Preliminaries and Notations
Online A/B experiments are widely used to compare customer responses between the existing feature A and the new feature B, guiding online companies’ feature launch decisions. To reduce the noise and established the causal relationship in online experiments, it is standard to track customers’ outcomes (e.g. ad clicks) after a customer is triggered based on some event (e.g., displaying ads).
Let the duration of the experiment be defined in terms of a certain number of weeks, which helps mitigate the influence of daily and weekend fluctuations. To maintain the simplicity, we assume the experiments have the same duration $t$ and omit the time subscript $t$ in notations. Denote $n_{g}$ as the total number of customers who are triggered up to week $t$ in group $g\in\{T,C\}$ (T for the treatment and C for control). Denote sets of those customers who are triggered up to week $t$ as $I_{g}$. Denote $Y_{i}$ as the observed outcome for customer $i$ up to week $t$, and $\bar{Y}_{g}=\sum_{i\in I_{g}}\frac{1}{n_{g}}Y_{i}$ is the sample mean of responses for customers in group $g$. Assuming that $Y_{i}$ are independent and identically distributed (i.i.d.) with mean $u_{g}$ and finite variance $\sigma^{2}_{g}$, as $n_{.}\rightarrow\infty$, we can apply the central limit theorem (CLT) (Casella and Berger 2021):
@@
$H_{0}:\mu\leq 0\quad H_{1}:\mu>0.$
@@
where $\alpha$ is Type I error. $\delta$ is AES. $z_{.}$ represents the standard normal quantile. $\Phi(.)$ stands for the cumulative distribution function of the standard normal distribution. To achieve a desired power of $b$ for the alternative hypothesis, considering for a given effect size $\delta>0$ and an allocation rate of $n_{T}/n_{C}=p$, the minimum sample size $n_{T},n_{C}$ are selected to satisfying $power(\delta,\sigma_{.},n_{.},\alpha)\geq b$. The required minimum duration can be determined through sample size prediction (Richardson et al. 2022).
Standardized effect sizes, such as Cohen’s d (Cohen 2013), are often used in meta-analysis across different experiments, serving to standardize variations in ATE magnitudes. If the estimators of standardized effect size follow an asymptotic normal distribution, all models presented in this paper can seamlessly accommodate the use of standardized effect sizes. Consequently, a dedicated discussion on this topic is omitted in this paper.
Pooled Effect Size
@@
The following model is used to estimate the AES (Du and Wang 2016):
@@
In an online company setting, not every experiment yields a measurable effect size. This stems from the nature of online experiments with their short cycles, which was designed to encourage the exploration of innovative ideas. However, the drawback is that a significant number of ideas may not achieve statistical significance. We categorize experiments into three groups: 1) true positive, 2) true negative, and 3) flat experiments which have no significant effect. Using $d_{i}$ from all experiments to train the model (Pooled Effect Size) will lead to an underestimation of AES, consequently elongating the required duration. The guideline is to train the model (Pooled Effect Size) using experiments where we know there is an underlying true positive effect. While it might seem intuitive to use hypothesis testing to categorize experiments into significant positive, negative or flat ones, the existence of Type I and Type II errors poses a challenge in precisely determining which experiments truly have such effects.
Three-Layer Heteroscedastic GMM
@@
(c) The simulated observed effect sizes, with three clusters corresponding to positive, flat, and negative.
@@
In the previous section, we highlighted that the main challenge of the state-of-the-art method lies in the accurately identifying experiments with true positive effect. It is natural to consider the categorization of experiments as latent variables and use the Gaussian Mixture Model (GMM).
Motivation. We first illustrate our motivation of using GMM in Figure 1. Figure 1(a) presents the distribution of the observed effect sizes for certain outcome of interest over 3,300 experiment run within Amazon. Although at the first glance, this graph seems to support a normal distribution, we argue that it indeed illustrates the challenge of this problem. In Figure 1(b), we simulate the true effect sizes following a two-Layer GMM, and in Figure 1(c) we increases random errors to each of them to generate the observed effect size, which looks like a single mode Gaussian distribution. Therefore, under the assumption that there exists latent clusters of positive, flat and negative experiments, it is actually infeasible to identify those positive ones with simple rules; instead, a more principled statistical model as GMM should be used. The two-layer GMM didn’t account for the required heterogeneity in our setting. Thus, we propose a three-layer heteroscedastic GMM.
@@
Latent variable $\bm{z}=\{z_{1},z_{2},\ldots,z_{m}\}$, $z_{i}\in\{1,\ldots,K\}$ represent the index class sampled from the categorical distribution parameterized by $\bm{\pi}=(\pi_{1},\ldots,\pi_{K})$. $\mu_{k},\tau^{2}_{k}$ and $\pi_{k}$ correspond to the mean, variance, and weight of the $k$-th Gaussian component. Denote $\bm{\theta}=\{\mu_{1},\ldots,\mu_{K},\tau_{1},\ldots,\tau_{K}\}$. We have
@@
where $\sigma^{2}_{i}$ are known and can be estimated using observed outcomes of the experiment $i$, e.g. Equation (1). We can further pre-specify the means of k = 2 cluster as zero, i.e. $\mu_{2}=0$.
@@
$=$ $\sum_{i=1}^{m}\mathbb{E}_{z|d}\bigg{[}log\bigg{(}p_{i}(d_{i},z_{i}|\bm{\theta},\bm{\pi})\bigg{)}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}$
$=$ $\sum_{i=1}^{m}\sum_{k=1}^{K}\mathbb{E}_{z|d}\bigg{[}{\mathbb{I}[z_{i}=k]}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}$
$\cdot\bigg{[}log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))\bigg{]}$
@@
The EM algorithm is summarized in Algorithm 1. By setting $K=3$ and $\mu_{2}=0$, the mean for positive components $\mu_{1}$ serves as the estimate of AES. To avoid getting trapped in stationary point(Wu 1983), we repeat the EM algorithm multiple times with randomly initialized starting points.
Singularity issues in Gaussian mixture model. When the covariance matrix is singular, the variances becomes zero, it leads to a spiky Gaussian component that ”collapses” into a single point. There are several papers discussing how to address this issue. The first approach involves setting a lower bound on variance (Hathaway 1985). The second approach introduces a penalty term in the log-likelihood function to prevent the variance of a specific component from becoming too small (Chen, Tan, and Zhang 2008; Ridolfi and Idier 2001). In addition, Chen, Tan, and Zhang (2008); Chen (2017) introduce the conditions and provide proofs for the consistency of the MLE under finite Gaussian mixture model. Using the second approach, the marginal log-likelihood takes the following form:
$log(f(\bm{d}|\bm{\theta},\bm{\pi}))+log(p_{m}(\bm{\tau^{2}}))=\sum_{i=1}^{m}\sum_{k=1}^{K}[log(\pi_{k})$
$+log(f_{k}(d_{i}|\bm{\theta}))]-\frac{1}{m}\sum_{k=1}^{K}(\frac{1}{\tau^{2}_{k}}+log(\tau^{2}_{k})),$ (10)
@@
Utility Theory: Bayesian Optimal Effect Size
In the previous section, we proposed a modeling approach that focuses on estimating the mean effect size. Such an approach is easy to explain to experimentation service users. However, in many real-world scenarios, users prioritize identifying the most suitable parameters for achieving maximum gains rather than just maximizing the accuracy of estimation. Motivated by the real needs, we propose a utility-maximization framework as our second approach.
@@
$\operatorname*{arg\;max}_{\mu>0}\;\;\sum_{i=1}^{m}{\mathbb{E}}_{\delta_{i}\sim\mathcal{N}(d^{\prime}_{i},\sigma_{i}^{2})}\Bigg{[}\underbrace{-c_{i}\cdot(T_{i}(\mu)-1)}_{\text{Opportunity Cost}}$
$+\underbrace{\big{(}\delta_{i}\!\cdot\!N_{i,1:T_{i}(\mu),T}\big{)}}_{\mathclap{\begin{subarray}{c}\text{Impact during the experiment}\\ \text{(relative to control)}\end{subarray}}}+\underbrace{u_{2}(\delta_{i},H,T_{i}(\mu))\mathbb{I}(\pi(\bm{Y}_{T_{i}(\mu),i})=1)}_{\text{Launch impact (relative to control)}}\Bigg{]}$
$=\operatorname*{arg\;max}_{\mu>0}\sum_{i=1}^{m}\Bigg{[}\underbrace{-c_{i}\cdot(T_{i}(\mu)-1)}_{\text{Opportunity Cost}}$
$+\underbrace{\big{(}d^{\prime}_{i}\!\cdot\!N_{i,1:T_{i}(\mu),T}\big{)}}_{\mathclap{\begin{subarray}{c}\text{Impact during the experiment}\\ \text{(relative to control)}\end{subarray}}}+\underbrace{u_{2}(d^{\prime}_{i},H,T_{i}(\mu))\mathbb{I}(\pi(\bm{Y}_{T_{i}(\mu),i})=1)}_{\text{Launch impact (relative to control)}}\Bigg{]}$
@@
  • $N_{i,1:t,g}$ denotes the total number of customers in group $g\in\{T,C\}$ observed up to week $t$ for experiment $i$.
@@
  • $u_{2}(\delta,H,t)$ represents the impact of launching the treatment feature on the whole population at week $t$, given a pre-specified time horizon $H$. Typically, we study one year impact, i.e. $H=52$ weeks. One example of estimating one year launch impact is: $u_{2}(\delta,H,t)=\delta*(H-t)*(\sum_{g\in\{T,C\}}N_{i,1:t,g})$
@@
The equality in Equation (Utility Theory: Bayesian Optimal Effect Size) is due to the linearity of expectation when $u_{2}$ is a linear function in the ATE (which is true in our case).
Optimization. Equation (Utility Theory: Bayesian Optimal Effect Size) is a one-dimensional optimization problem, for which we can efficiently find a good solution. We use grid search in our prototype.
@@
Accuracy Comparison with Simulation
@@
Methods and Results. Table 1 compares the MSE and MAE for the pooled effect size (Pooled-MLE), standard two-layer GMM (Two-layer GMM) and the proposed three-layer heteroscedastic GMM (Three-layer GMM). For Pooled-MLE, the MLE of parameter $\mu_{0}$ is computed using the model in Equation (Pooled Effect Size) on all positive observed effect sizes $d_{i}$. Two-layer GMM uses Gaussian Mixture in sklearn package (Pedregosa et al. 2011) with $K=3$. Three-layer GMM, the proposed method, uses the proposed EM-Algorithm 1 with the penalty term (Three-Layer Heteroscedastic GMM), $K=3$ and $\mu_{2}=0$. Each simulation uses $10$ different starting points to avoid local optimum (One of these starting points is initialized using the k-means clustering). Both Two-layer GMM and Three-layer GMM uses the estimated mean for the positive components as AES and $tolerance=10^{-3}$. Table 1 shows that Three-layer GMM performs slightly better than Two-layer GMM. The p-value from a two-sample t-test is 0.036, indicating the Three-layer GMM has significantly better accuracy. As expected, Pooled-MLE is underestimated. GMM outperforms Pooled-MLE notably in cases where we lack information about whether the experiment’s effect is a true positive or not. The Boxplot in Figure 3 illustrates a similar conclusion.
@@
Meta-Analysis with Real Experiments
@@
Dataset and setup. We collect a dataset of 3,300 historical experiments conducted in the past two years within Amazon, each having a duration of 4 weeks. $d_{i}$ represents the observed standardized effect size derived from a specific standardization formulation used in the company at week 4. Given the absence of the ground-truth effect size, we adopt a heuristic yet easy-to-explain approach that uses the observed effect size at the end of the 4 weeks as the empirical ground-truth effect size.
We set the maximum duration as $4$ weeks for the analysis. For each estimated AES obtained from different methods, we plug in them, along with the observed sample size and sample variance (or their predicted versions as described in Richardson et al. (2022)), into power Equation (2) to calculate the statistical power at week $i=1,2,3,4$. We define the recommended duration as the minimum number of weeks ($\leq 4$) required to attain a power of 80%. In Frequentist setting, we use the one-sided two-group Welch’s t-test (Welch 1947) as decision policy $\pi(.)$. This implies that the decision policy $\pi(.)=1$ for launch if the p-value is less than the significance level of $0.05$ and the ATE is positive.
@@
Table 3: Simulated data analysis results (Ground truth is 1).
Methods. We compare the proposed estimated AESs from Three-Layer GMM and utility theory-based optimal effect size (Utility-maximization) with those from two-layer GMM and Pooled-MLE. The settings for Two-layer GMM, Three-layer GMM and Pooled_MLE are the same as the previous section. Utility-maximization uses grid-search to find the optimal effect size within the set $\{0.02\%,0.04\%,\ldots,2\%\}$.
@@
    The average utility in Equation (Utility Theory: Bayesian Optimal Effect Size) and its three components, including the opportunity cost, the impact during the experiment, and the launch impact.